OpenStudy (anonymous):

Help!!!! (with part B, I need help with that) An investment fund has increased in value by a total of 21% over the last 2 years. a) A man invested $1000 in the fund 2 years ago. What is it worth now? b) Calculate the yearly rate of interest assuming it was: i) Compound interest ii) Simple interest

OpenStudy (gorv):

okay..let the fund was =x initially

OpenStudy (gorv):

now first by simple interest

OpenStudy (gorv):

\[interest=\frac{ amount *rate *time }{ 100}\]

OpenStudy (gorv):

it is given fund increase by 21 % after 2 year so after 2 year fund = x+x*21/100

OpenStudy (gorv):

okk??

OpenStudy (anonymous):

Are you mad?

OpenStudy (gorv):

why???

OpenStudy (anonymous):

Because you're saying 'okk??' It made me think you're angry or something. Nevermind

OpenStudy (gorv):

lol i mean u got me ????

OpenStudy (gorv):

u got what i told upto here??

OpenStudy (anonymous):

Yeah I understand :)

OpenStudy (gorv):

that what i mean ??

OpenStudy (anonymous):

Yes :3

OpenStudy (gorv):

so after 2 year fund = x+x*21/100 final amount = x+x*21/100 initail = x can u tell me what amount is add to x in last 2 year

OpenStudy (anonymous):

I'm getting confused....

OpenStudy (gorv):

u initailly had x now u had x+x*21/100 what amount is added to x

OpenStudy (anonymous):

so the interest is 1000*21%*2 years right?

OpenStudy (gorv):

like earlier u had 10 dollar now u have 25 what amount added to initia

OpenStudy (anonymous):

15

OpenStudy (gorv):

smilarly do for x

OpenStudy (gorv):

tell me what increment happend in x in 2 year

OpenStudy (anonymous):

21% ?

OpenStudy (gorv):

how

OpenStudy (gorv):

x+x*21/100-x=??

OpenStudy (gorv):

\[x+x*\frac{ 21 }{ 100 }-x\]

OpenStudy (gorv):

now can u tell buddy :??

OpenStudy (anonymous):

I don't know... 21?

OpenStudy (gorv):

\[x-x+\frac{ 21*x }{ 100 }\]

OpenStudy (gorv):

x will get cancelled and we left with \[\frac{ 21*x }{ 100 }\]

OpenStudy (anonymous):

I see

OpenStudy (gorv):

so this is our interest

OpenStudy (anonymous):

210?

OpenStudy (anonymous):

Is 210 the interest?

OpenStudy (gorv):

what??

OpenStudy (gorv):

how u got it ??

OpenStudy (gorv):

21*x/100 i told u na

OpenStudy (gorv):

\[\frac{ 21x }{ 100 }=\frac{ x*r*2 }{ 100 }\]

OpenStudy (gorv):

r=rate t=2 year principle= x

OpenStudy (anonymous):

Yes

OpenStudy (gorv):

divide both side by x

OpenStudy (anonymous):

So x gets cancelled?

OpenStudy (gorv):

21/100=(r*2)/100

OpenStudy (gorv):

yeah now multiply by 100

OpenStudy (anonymous):

So the answer is??

OpenStudy (gorv):

after multiplying by 100 what we will get

OpenStudy (anonymous):

You tell me, I don't know

OpenStudy (gorv):

u only need ans :(

OpenStudy (anonymous):

But I don't know :(